OpenStudy (anonymous):

Complete the table of ordered pairs for the equation 5x-7y=-35. x / y 0 0 -2 -5

OpenStudy (aravindg):

35/7

OpenStudy (aravindg):

-7

OpenStudy (aravindg):

25/7

OpenStudy (aravindg):

-70/5

OpenStudy (aravindg):

the process is simple put the respective value of x or y and then solve for the other variable

OpenStudy (aravindg):

i will show u hw i gt first one

OpenStudy (aravindg):

5x-7y=-35 5(0)-7y=-35 -7y=-35 y=35/7

OpenStudy (anonymous):

Oh I see how you got there now! Thank you for helping me out ;)